# Uplink-Downlink Duality for Beamforming in Integrated Sensing and Communications

### v1

@@ I Introduction
To quantify the sensing performance, we utilize the Cramér-Rao bound (CRB) as a metric for sensing. The optimization of CRB is complicated due to the fact that the structure of the CRB involves a matrix inverse (e.g., see [6, 7, 8] for the scalar case where the CRB is a fraction). One of the main contributions of this paper is that, under certain conditions, the minimization of CRB can be transformed into a sequence of subproblems, each involving maximizing power in certain sensing directions of interest. This essentially reduces the ISAC problem to downlink beamforming problems that resemble those treated in classical works in MIMO communications [9, 10, 11, 12, 13, 14]. These subproblems are indexed by a set of auxiliary variables, which can be optimized in an outer loop.
@@ I-C Paper Organization and Notation
This paper uses lowercase, lowercase boldface, and uppercase boldface letters to denote scalars, vectors, and matrices, respectively. We use $(\cdot)^{\textsf{T}}$, $(\cdot)^{\textsf{H}}$, $(\cdot)^{-1}$, and $(\cdot)^{\dagger}$ to represent the transpose, Hermitian, inverse, and pseudo-inverse of a matrix. We use $\Re\{\cdot\}$, $\Im\{\cdot\}$ to denote the real and imaginary parts of a complex number, $\overline{(\cdot)}$ as the complex conjugate, and $\imath$ as $\sqrt{-1}$. For vectors, $\|\cdot\|$ denotes the Euclidean norm. The matrix operators $\operatorname*{Tr}(\cdot)$ and $\rho_{\text{max}}(\cdot)$ denote the trace and spectral radius. Note that for a PSD matrix, the spectral radius is its maximum eigenvalue. We use $\mathbf{I}$ to denote an identity matrix of appropriate dimensions with its $\ell$-th column denoted by $\mathbf{e}_{\ell}$. We use $\mathbf{1}$ to denote the vector of all ones and $\mathbf{0}$ a matrix (or vector) of all zeros with appropriate dimension. We use $\operatorname*{diag}(d_{1},\ldots,d_{K})$ to denote a diagonal matrix with diagonal elements $d_{1},\ldots,d_{K}$. The space of $n\times m$ matrices with real (complex) entries is denoted by $\mathbb{R}^{n\times m}$ ($\mathbb{C}^{n\times m}$). We use $\mathbb{E}[\cdot]$ to denote the expectation of a random variable. Finally, $\mathcal{CN}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ represents the circularly-symmetric complex Gaussian vector with mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$.
@@ II-B Sensing Model and Performance Metric
The sensing model adopted in this paper assumes that the BS can perform self-interference mitigation [34], i.e., the leakage from the transmit array to the received antennas can be cancelled. Consequently, the baseband signal signal $\mathbf{y}_{\text{s}}[t]\in\mathbb{C}^{N_{\text{R}}}$ received at the BS during the $t$-th symbol period can be written as
@@
where $\mathbf{G}^{\left(\boldsymbol{\eta}\right)}\in\mathbb{C}^{N_{\text{R}}\times N_{\text{T}}}$ is the sensing channel matrix that models the “round-trip-return” channel between the BS transmit and receive antenna arrays, and $\mathbf{n}_{\text{s}}[t]\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I})$ is the additive noise at the receiver.
@@
$\mathbf{T}_{\mathbf{V}}]_{ij}=\mathbb{E}_{\boldsymbol{\eta}}\left[\frac{T}{\sigma^{2}}\operatorname*{Tr}\left(\ddot{\mathbf{G}}_{ij}^{(\boldsymbol{\eta})}\mathbf{V}\mathbf{V}^{\textsf{H}}\right)\right],$ (8)
@@
Asymptotically, either as $\sigma^{2}\rightarrow 0$ or as $T\rightarrow\infty$, the BCRB provides a tight bound on the mean-squared error (MSE) matrix associated with the optimal (i.e., the minimum mean-squared error (MMSE)) estimator.
@@
We remark that the BCRB considered in this work coincides with the notion in Van Trees [35] that treats $(\mathbf{X},\mathbf{Y}_{\text{s}})$ as random observations revealed to the estimator. In the literature, tighter bounds on the MSE are also known, e.g., the Weiss-Weinstein bound [37] and the Ziv-Zakai bound [38]. Furthermore, even within the CRB family, the so-called Miller-Chang bound [39] is reported to be tighter in some scenarios. Instead of setting the weights corresponding to the nuisance parameters to be zero, the Miller-Chang technique [39] takes statistical average over the nuisance parameters in the computation of CRB. However, these potentially tighter bounds are also more difficult to compute and to optimize. In this paper, we restrict attention to the BCRB as given by the BFIM in (6)-(8).
The main advantage of using the BCRB as the sensing metric is that the BFIM has a closed-form that is relatively simple to compute as compared to the exact MSE calculation. Furthermore, it abstracts the receiver design since the bound holds for all practical estimators. Finally, we note that unlike the classical CRB used in previous works [6, 15], the BCRB does not depend on the actual value of the unknown parameter $\boldsymbol{\eta}$ and only the prior $f(\boldsymbol{\eta})$.
@@ II-C Communication Model and Performance Metric
where $\mathbf{h}_{k}\in\mathbb{C}^{N_{\text{T}}}$ is the communication channel of the $k$-th user, and $n_{k}[t]\sim\mathcal{CN}\left(0,\sigma^{2}\right)$ is the additive white Gaussian noise. In this paper, we make the assumption that the communication channel $\mathbf{H}\triangleq\left[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ can be perfectly estimated within each coherence interval (e.g., using uplink pilots and assuming channel reciprocity), and is known perfectly at the BS.
@@
$\text{SINR}^{\text{DL}}_{k}(\mathbf{V})\triangleq\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},$ (11)
where $\gamma_{k}$ denote the SINR threshold for the $k$-th user. Throughout this paper, we assume that these SINR thresholds are strictly feasible.
@@ II-D Problem Formulation
$\mathrm{subject\ to}$~~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k,$ (12b)
@@
Throughout this paper, we assume that the ISAC problem with the communication channel $\mathbf{H}$, the sensing channel $\mathbf{G}^{(\boldsymbol{\eta})}$, the prior $f(\boldsymbol{\eta})$, and power constraint $P$, has the following property. Let $\mu\in\mathbb{R}$ denote the trace of weighted BCRB, and let $\boldsymbol{\gamma}\in\mathbb{R}^{K}$ denote a vector of downlink communication SINR thresholds. We assume that whenever $(\mu,\boldsymbol{\gamma})$ is achievable using the extended model (2), i.e., there exists an extended matrix $\mathbf{V}_{\text{ex}}=\left[\begin{matrix}\mathbf{V}&\mathbf{V}_{\text{s}}\end{matrix}\right]$ with a communication beamforming matrix $\mathbf{V}=\left[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ and a sensing beamformer matrix $\mathbf{V}_{\text{s}}\in\mathbf{C}^{N_{\text{T}}\times N_{\text{T}}}$ achieving
@@
$\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\mathbf{h}_{k}^{\mathsf{H}}\mathbf{V}_{\text{s}}\mathbf{V}_{\text{s}}^{\mathsf{H}}\mathbf{h}_{k}+\sigma^{2}}\geq\gamma_{k},\quad\forall k,$ (14)
@@
The above assumption may seem restrictive at first glance, but recent studies have established several sufficient conditions under which Assumption 1 holds. For instance, as demonstrated in [36], when the number of users $K$ is large as compared to the number of parameters to be estimated $L$, more precisely, when
$K\geq\tfrac{L(L+1)}{4},$ (15)
@@ II-E Example
$\mathbf{a}_{\text{T}}(\theta)=\left[e^{\imath\pi 0\sin(\theta)},\ldots,e^{\imath\pi\left(N_{\text{T}}-1\right)\sin(\theta)}\right]^{\textsf{T}}.$
@@
Now, assume a Gaussian prior for $\alpha$ and $\theta$ with variances $\tfrac{\sigma_{\alpha}^{2}}{2T}\sigma^{2}$ and $\tfrac{\sigma_{\theta}^{2}}{2T}\sigma^{2}$, respectively, and arbitrary means. In this case, the BFIM in (6) can be shown to be
$\mathbf{J}_{\mathbf{V}}=\frac{2T}{\sigma^{2}}\left[\begin{matrix}h_{\alpha}(\mathbf{V})+\frac{1}{\sigma_{\alpha}^{2}}&0&\Re\{h_{\theta\alpha}(\mathbf{V})\}\\ 0&h_{\alpha}(\mathbf{V})+\frac{1}{\sigma_{\alpha}^{2}}&\Im\{h_{\theta\alpha}(\mathbf{V})\}\\ \Re\{h_{\theta\alpha}(\mathbf{V})\}&\Im\{h_{\theta\alpha}(\mathbf{V})\}&h_{\theta}(\mathbf{V})+\frac{1}{\sigma_{\theta}^{2}}\end{matrix}\right]$ (18)
@@
$\left[\mathbf{J}^{-1}\right]_{3,3}=(s-\mathbf{r}^{\mathsf{H}}\mathbf{S}\mathbf{r})^{-1},$ (22)
@@
Thus, for an ISAC system serving at least two users, while estimating an AoA at the same time, Assumption 1 holds, i.e., using only communication beams can already achieve the performance of the extended model with extra sensing beams.
@@ II-F Existing Method and Proposed Solution Strategy
For minimizing the sum power (i.e., when $\mathbf{\tilde{Q}}=\mathbf{I}$), a duality notion has been developed in [10] that allows the downlink problem (26) to be transformed to a virtual uplink problem, a result now commonly referred to as uplink-downlink duality [10, 11, 26, 27, 12]. The key advantage of such a transformation is that the uplink problem is more computationally efficient to solve than its downlink counterpart. The duality relationship for arbitrary $\mathbf{\tilde{Q}}\succcurlyeq 0$ is developed in [12]. This paper further develops the uplink-downlink duality theory for the BCRB minimization problem. We show how the sensing objective in (12) can be transformed, and how the classical uplink-downlink duality theory can be extended for the transformed problem.
@@ III-A Semidefinite Relaxation
$\mathrm{subject\ to}$~ $\frac{1}{\gamma_{k}}\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}_{k}\mathbf{h}_{k}-\sum_{i\neq k}\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}_{i}\mathbf{h}_{k}\geq\sigma^{2}$ (27b)
@@
where $\mathbf{J}_{\mathbf{R}}$ denotes, with a slight abuse of notation, the BFIM matrix expressed in $\mathbf{R}_{1},\ldots,\mathbf{R}_{k}$ (i.e., with $\mathbf{V}\mathbf{V}^{\textsf{H}}$ in (8) replaced by $\sum_{k}\mathbf{R}_{k}$), and the rank-one constraints on $\mathbf{R}_{k}$ are dropped.

### v2

@@ I Introduction
To quantify the sensing performance, we utilize the Cramér-Rao bound (CRB) as a metric for sensing. The optimization of CRB is not straightforward due to the fact that the structure of the CRB involves a matrix inverse (e.g., see [6, 7, 8] for the scalar case where the CRB is a fraction). One of the main contributions of this paper is that, under certain conditions, the minimization of CRB can be transformed into a sequence of subproblems, each involving maximizing power in certain sensing directions of interest. This essentially reduces the ISAC problem to downlink beamforming problems that resemble those treated in classical works in MIMO communications [9, 10, 11, 12, 13, 14]. These subproblems are indexed by a set of auxiliary variables, which can be optimized in an outer loop.
@@ I-C Paper Organization and Notation
This paper uses lowercase, lowercase boldface, and uppercase boldface letters to denote scalars, vectors, and matrices, respectively. We use $(\cdot)^{\textsf{T}}$, $(\cdot)^{\textsf{H}}$, $(\cdot)^{-1}$, and $(\cdot)^{\dagger}$ to represent the transpose, Hermitian, inverse, and pseudo-inverse of a matrix. We use $\Re\{\cdot\}$, $\Im\{\cdot\}$ to denote the real and imaginary parts of a complex number, $\overline{(\cdot)}$ as the complex conjugate, and $\imath$ as $\sqrt{-1}$. For vectors, $\|\cdot\|$ denotes the Euclidean norm. The matrix operators $\operatorname*{Tr}(\cdot)$ and $\rho_{\text{max}}(\cdot)$ denote the trace and spectral radius. Note that for a PSD matrix, the spectral radius is its maximum eigenvalue. We use $\mathbf{I}$ to denote an identity matrix of appropriate dimensions with its $\ell$-th column denoted by $\mathbf{e}_{\ell}$. We use $\mathbf{1}$ to denote the vector of all ones and $\mathbf{0}$ a matrix (or vector) of all zeros with appropriate dimension. We use $\operatorname*{diag}(d_{1},\ldots,d_{K})$ to denote a diagonal matrix with diagonal elements $d_{1},\ldots,d_{K}$. The space of $n\times m$ matrices with real (complex) entries is denoted by $\mathbb{R}^{n\times m}$ ($\mathbb{C}^{n\times m}$). The set of vectors in $\mathbb{R}^{K}$ with nonnegative (positive) entries is denoted by $\mathbb{R}^{K}_{+}$ ($\mathbb{R}^{K}_{++}$). We use $\mathbb{E}[\cdot]$ to denote the expectation of a random variable. Finally, $\mathcal{CN}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ represents the distribution of circularly symmetric complex Gaussian vectors with mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$.
@@ II-B Sensing Model and Performance Metric
The sensing model adopted in this paper assumes that the BS can perform self-interference mitigation [34], i.e., the leakage from the transmit array to the received antennas can be cancelled. Consequently, the baseband signal signal $\mathbf{y}_{\text{s}}[t]\in\mathbb{C}^{N_{\text{R}}}$ received at the BS for the $t$-th symbol (after removing the effect of channel delay) can be written as
@@
where $\mathbf{G}^{\left(\boldsymbol{\eta}\right)}\in\mathbb{C}^{N_{\text{R}}\times N_{\text{T}}}$ is the sensing channel matrix that models the “round-trip-return” channel between the BS transmit and receive antenna arrays, and $\mathbf{n}_{\text{s}}[t]\sim\mathcal{CN}(0,\sigma_{\text{s}}^{2}\mathbf{I})$ is the additive noise at the receiver.
@@
$\mathbf{T}_{\mathbf{V}}]_{ij}=\mathbb{E}_{\boldsymbol{\eta}}\left[\frac{T}{\sigma_{\text{s}}^{2}}\operatorname*{Tr}\left(\ddot{\mathbf{G}}_{ij}^{(\boldsymbol{\eta})}\mathbf{V}\mathbf{V}^{\textsf{H}}\right)\right],$ (8)
@@
Asymptotically, either as $\sigma_{\text{s}}^{2}\rightarrow 0$ or as $T\rightarrow\infty$, the BCRB provides a tight bound on the mean-squared error (MSE) matrix associated with the optimal (i.e., the minimum mean-squared error (MMSE)) estimator.
@@
We remark that the BCRB considered in this work coincides with the notion in standard textbook in estimation (e.g., [35]) that treats $(\mathbf{X},\mathbf{Y}_{\text{s}})$ as random observations revealed to the estimator. In the literature, tighter bounds on the MSE are also known, e.g., the Weiss-Weinstein bound [37] and the Ziv-Zakai bound [38]. Furthermore, even within the CRB family, the so-called Miller-Chang bound [39] is reported to be tighter in some scenarios. Instead of setting the weights corresponding to the nuisance parameters to be zero, the Miller-Chang technique [39] takes statistical average over the nuisance parameters in the computation of CRB. However, these potentially tighter bounds are also more difficult to compute and to optimize in most cases. In this paper, we restrict attention to the BCRB as given by the BFIM in (6)-(8).
The main advantage of using the BCRB as the sensing metric is that the BFIM has a closed-form that is relatively simple to compute as compared to the exact MSE calculation. Furthermore, it abstracts the receiver design since the bound holds for all practical estimators. Further, we note that unlike the classical CRB used in previous works [6, 15], the BCRB does not depend on the actual value of the unknown parameter $\boldsymbol{\eta}$ and only the prior $f(\boldsymbol{\eta})$.
We remark that the preceding formulation can be applied to the setting of active sensing [40], where sensing occurs adaptively across multiple stages of the coherence interval, by updating the prior from one stage to the next as the BS learns more about the parameter [16]. This can improve the sensing performance, but at the cost of increased computational complexity of optimizing the sensing strategy in each stage.
@@ II-C Communication Model and Performance Metric
where $\mathbf{h}_{k}\in\mathbb{C}^{N_{\text{T}}}$ is the communication channel of the $k$-th user, and $n_{k}[t]\sim\mathcal{CN}\left(0,\sigma_{\text{c}}^{2}\right)$ is the additive white Gaussian noise. In this paper, we make the assumption that the communication channel $\mathbf{H}\triangleq\left[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ can be perfectly estimated within each coherence interval (e.g., using uplink pilots and assuming channel reciprocity), and is known perfectly at the BS.
@@
$\text{SINR}^{\text{DL}}_{k}(\mathbf{V})\triangleq\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},$ (11)
where $\gamma_{k}$ denote the SINR threshold for the $k$-th user. Throughout this paper, we assume that $\gamma_{k}$ is positive for all $k$ (i.e., the users have positive rate constraints), and that the SINR constraints are strictly feasible.
@@ II-D Problem Formulation
$\mathrm{subject\ to}$~~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},\quad\forall k,$ (12b)
@@
Throughout this paper, we assume that the ISAC problem with the communication channel $\mathbf{H}$, the sensing channel $\mathbf{G}^{(\boldsymbol{\eta})}$, the prior $f(\boldsymbol{\eta})$, and power constraint $P$, has the following property. Let $\mu\in\mathbb{R}_{+}$ denote the trace of weighted BCRB, and let $\boldsymbol{\gamma}\in\mathbb{R}^{K}_{++}$ denote a positive vector of downlink communication SINR thresholds. We assume that whenever $(\mu,\boldsymbol{\gamma})$ is achievable using the extended model (2), i.e., there exists an extended matrix $\mathbf{V}_{\text{ex}}=\left[\begin{matrix}\mathbf{V}&\mathbf{V}_{\text{s}}\end{matrix}\right]$ with a communication beamforming matrix $\mathbf{V}=\left[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ and a sensing beamformer matrix $\mathbf{V}_{\text{s}}\in\mathbf{C}^{N_{\text{T}}\times N_{\text{T}}}$ achieving
@@
$\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\mathbf{h}_{k}^{\mathsf{H}}\mathbf{V}_{\text{s}}\mathbf{V}_{\text{s}}^{\mathsf{H}}\mathbf{h}_{k}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},\quad\forall k,$ (14)
@@
The preceding assumption may seem restrictive at first glance, but recent studies have established several sufficient conditions under which Assumption 1 holds. For instance, as demonstrated in [36], when the number of users $K$ is large as compared to the number of parameters to be estimated $L$, more precisely, when
$K\geq\frac{L(L+1)}{4},$ (15)
@@ II-E Example
$\mathbf{a}_{\text{T}}(\theta)=\frac{1}{\sqrt{N_{\text{T}}}}\left[e^{\imath\pi 0\sin(\theta)},\ldots,e^{\imath\pi\left(N_{\text{T}}-1\right)\sin(\theta)}\right]^{\textsf{T}}.$
@@
Now, assuming a Gaussian prior for $\alpha$ and $\theta$ with variances $\tfrac{\sigma_{\alpha}^{2}}{2T}\sigma_{\text{s}}^{2}$ and $\tfrac{\sigma_{\theta}^{2}}{2T}\sigma_{\text{s}}^{2}$, respectively, and arbitrary means, the BFIM in (6) can be shown to be
$\mathbf{J}_{\mathbf{V}}=\frac{2T}{\sigma_{\text{s}}^{2}}\left[\begin{matrix}h_{\alpha}(\mathbf{V})+\frac{1}{\sigma_{\alpha}^{2}}&0&\Re\{h_{\theta\alpha}(\mathbf{V})\}\\ 0&h_{\alpha}(\mathbf{V})+\frac{1}{\sigma_{\alpha}^{2}}&\Im\{h_{\theta\alpha}(\mathbf{V})\}\\ \Re\{h_{\theta\alpha}(\mathbf{V})\}&\Im\{h_{\theta\alpha}(\mathbf{V})\}&h_{\theta}(\mathbf{V})+\frac{1}{\sigma_{\theta}^{2}}\end{matrix}\right]$ (18)
@@
$\left[\mathbf{J}^{-1}\right]_{3,3}=(s-\mathbf{r}^{\mathsf{H}}\mathbf{S}^{-1}\mathbf{r})^{-1},$ (22)
@@
Thus, for an ISAC system serving at least two communication users, while estimating an AoA at the same time, Assumption 1 holds, i.e., using only communication beams can already achieve the performance of the extended model with extra sensing beams.
@@ II-F Existing Method and Proposed Solution Strategy
For minimizing the sum power (i.e., when $\mathbf{\tilde{Q}}=\mathbf{I}$), a duality notion has been developed in [10] that allows the downlink problem (26) to be transformed to a virtual uplink problem, a result now commonly referred to as uplink-downlink duality [10, 11, 26, 27, 12]. The key advantage of such a transformation is that the uplink problem is more computationally efficient to solve than its downlink counterpart. The duality relationship for arbitrary $\mathbf{\tilde{Q}}\succcurlyeq 0$ is developed in [12]. This paper further develops the uplink-downlink duality theory for the BCRB minimization problem, for which the effective $\mathbf{\tilde{Q}}$ may not necessarily be PSD. We show how the sensing objective in (12) can be transformed, and how the classical uplink-downlink duality theory can be extended for the transformed problem.
@@ III-A Semidefinite Relaxation
$\mathrm{subject\ to}$~ $\frac{1}{\gamma_{k}}\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}_{k}\mathbf{h}_{k}-\sum_{i\neq k}\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}_{i}\mathbf{h}_{k}\geq\sigma_{\text{c}}^{2}$ (27b)
@@
where $\mathbf{J}_{\mathbf{R}}$ denotes, with a slight abuse of notation, the BFIM matrix expressed in $\mathbf{R}_{1},\ldots,\mathbf{R}_{k}$ (i.e., with $\mathbf{V}\mathbf{V}^{\textsf{H}}$ in (8) replaced by $\sum_{k}\mathbf{R}_{k}$), and the rank-one constraints on $\mathbf{R}_{k}$ are dropped. Note that $\tfrac{1}{\gamma_{k}}$ is finite, since $\gamma_{k}>0$ for all $k$, so the problem is well defined.